OpenStudy (anonymous):

Complete each statement. If 4a=45, then a/5= ____ and a/9= ____ . if 10k=25, then 2k=___ and k/5=___. if m+p/p = 12/5, m/p=___, and p/m___.

hartnn (hartnn):

4a = 45 a =45/4 then a/5 =... ?

OpenStudy (anonymous):

where did 5 came from?

hartnn (hartnn):

it was in the question!

hartnn (hartnn):

If 4a=45, then a/5= ____

OpenStudy (anonymous):

oh sorryy.. yes, that's the question. but I could not figure it out.

hartnn (hartnn):

a =45/4 did u get this ?

OpenStudy (anonymous):

9 x 5 = 45 so.. a/5= a/9 ?

OpenStudy (anonymous):

or maybe 9/4 ?

hartnn (hartnn):

\(\huge \color{blue}{4a=45 \\a=\frac{45}{4}\\ \frac{a}{5}=\frac{45}{4*5}=\frac{9}{4}}\) yes, 9/4 is correct.

hartnn (hartnn):

what about a/9 =.. ?

OpenStudy (anonymous):

and a/9 would be 5/4

hartnn (hartnn):

\(\huge \color{blue}{4a=45 \\a=\frac{45}{4}\\ \frac{a}{9}=\frac{45}{4*9}=\frac{5}{4}}\) good!

hartnn (hartnn):

solve the 2nd one in same way.....10k =25 k=... ?

OpenStudy (anonymous):

how is it like? its 2k= __ ? k=25/10. =5*5/10 I am confused.

hartnn (hartnn):

you are going good. \(\huge \color{blue}{10k=25 \\k=\frac{25}{10}\\2k=\frac{2*5*5}{2*5}=...? }\)

OpenStudy (anonymous):

2k = 5 ? divide both sides by 5

hartnn (hartnn):

yes, u can do that also, and u get k = 5/2

hartnn (hartnn):

and 2k what was asked for =5 what about k/5=... ?

OpenStudy (anonymous):

k/5 = k/2

hartnn (hartnn):

how you getting k on right side also ??

hartnn (hartnn):

k/5 is just 1/2

OpenStudy (anonymous):

Wait.. if k= 25/10 2k = 50 / 10 2k = 5 so.. k/5 = 1/2 because you multiply 2 and 5 in both sides. Thanks!

hartnn (hartnn):

yes! good :)

hartnn (hartnn):

because you 'divide' 2 and 5 in both sides.

OpenStudy (anonymous):

why is 10 not multiplied twice? but 25 is multiplied.

hartnn (hartnn):

because '2' was multiplied to the 'numerator' of left side. so 2 will be multiplied only to numerator of right side, that is to 25 only.

OpenStudy (anonymous):

Ohh. Thankyou!

hartnn (hartnn):

:) tried 3rd ?

OpenStudy (anonymous):

yes. m/p would be 7/5 ? and p/m = 5/7 did I get it correctly??

hartnn (hartnn):

wow! absolutely correct \(\huge \checkmark\)

OpenStudy (anonymous):

Finally. Yey! :^_^